OpenStudy (ashley_fay):

a right triangle has a hypotenuse of length 17 and a leg of length 13. what is the area of the triangle

OpenStudy (stamp):



OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Area of triangle = (1/2)bh In this triangle, x is the base (b) and 13 is the height (h) With the Pythagoras theorem you can find x x^2 + 13^2 = 17^2